If given to questions A/2=B/3=C/4 then find value of a:b:c

Grade:12

3 Answers

VIBHU SHARAN DIXIT
6 years ago
firstly obtain the values of A and C in the terms of B viz:A=2B/3, B=B, C=4B/3
now by solving these terms the ratio obtained is 2/3:1:4/3
Sangeetha Prabakar
5 years ago
First : A/B = 2/3
Second: C/B = 4/3
So, A= 2; B = 3; C = 4
So, A:B:C = 2:3:4
Hope you understand this simple method. Let me know please. Happy to help.
Manda Snehashish Reddy
5 years ago
If the given condition is considered then sinA:sinB:sinC = 2:3:4
multiply with 2R in the sin ratios
2RsinA:2RsinB:2RsinC = 2:3:4
but, we know that a=2RsinA
 
Thus, a:b:c is also 2:3:4